PCS just paid a $3 dividend. Expected annual dividend growth is 3% per year from Year 1 to 3; 4% per year from Year 4 to 6 and 6% constant dividend. Required rate of return is 12%. What is the current share price of PCS? Price at Year 5?
Answer #1

D0 =3
D1 =3*(1+3%)
D2=3*(1+3%)^2
D3=3*(1+3%)^3
D4=3*(1+3%)^3*(1+4%)
D5=3*(1+3%)^3*(1+4%)^2
D6=3*(1+3%)^3*(1+4%)^3
D7 =3*(1+3%)^3*(1+4%)^3*(1+6%)

Terminal Value =D7/(Required Rate-Growth) =3*(1+3%)^3*(1+4%)^3*(1+6%)/(12%-6%) =65.14597

Current Price of share
=3*(1+3%)/(1+12%)+3*(1+3%)^2/(1+12%)^2+3*(1+3%)^3/(1+12%)^3+3*(1+3%)^3*(1+4%)/(1+12%)^4
+3*(1+3%)^3*(1+4%)^2/(1+12%)^5+3*(1+3%)^3*(1+4%)^3/(1+12%)^6+65.14597/(1+12%)^6=46.68

Price at year 5 =(Terminal Value+D6)/(1+r) =(65.14597+3*(1+3%)^3*(1+4%)^3)/(1+12%) =61.46


​​​​​​​